can you solve x with sin(x)=1/5 in the interval [0,2pi] algebraically...or without a graph? ...

Answer by stanbon(75887)

can you solve x with sin(x)=1/5 in the interval [0,2pi] algebraically...or without a graph?
--------------------------------------
x = arcsin(1/5) = 0.20 radians , or x = pi-0.2 = 2.94 radians
==================
Cheers,
Stan H.
==================